There are two right circular cylinders. The radius of the first cylinder is 4 centimeters, and its height is 5 centimeters. The radius of the second cylinder is 12 centimeters, and its height is also 5 centimeters. What is the ratio of the volume of the larger cylinder to the volume of the smaller cylinder
Answer by sudhanshu_kmr(1152)


first volume = pi* 4^2 * 5

second volume = pi* 12^2 * 5
second is larger, first is smaller cylinder.
ratio of volume = pi* 12^2 * 5/ pi* 4^2 * 5

= 144/16
= 9
ratio is : 9:1